# Affordances and Design Principles of The Political Left and Right

### v1

@@ 5. Discussion
Our findings continue to present the relevance of political ideologies as value frameworks in the context of technology design. We observed that affordances were used diversely between the particular leaning groups; we also observed that the mixed group seemed to adapt ideas that were present in right-leaning and left-leaning groups, speaking on their ability to develop ideas across sectors.
@@ 5.1. Design by the Left and the Right
Political science has various characterisations of the left and the right: citizens and experts identify this continuum and can position themselves into it (see e.g., Cochrane, 2015b). Using political party manifestos, scholars such as Cochrane (2015a), Jahn (2011) and Däubler and Benoit (2022) have identified characteristics of right-leaning parties to have – at least a subset of – ideals such as less state intervention or support for more expansive welfare programs characterise this tension (see Table 5, first row). As these are from political party manifestos, these do not directly map into the design of digital services but rather broadly characterise the division between the left and the right around the economy and the state.
@@
Moving from political manifestos to digital artifacts and design principles in them (see Table 5, second row row). Grön and Nelimarkka (2020) suggested that the left–right dimension appeared as a tension of paternalims and freedom. More extensively, they showed acceptable design approaches seemed to be rooted in the values of each party, for example, communality was preferred by a liberal-conservative and agrarian Centre Party.
We observe a similar ideological division in the role of market dynamics in controlling visibility. The right-leaning group suggested affordances encouraging popularity as a driver for content visibility, with an underlying idea of the platform as a marketplace for ideas competing for attention. Both left-leaning groups discouraged such an approach. Although traditional political party manifestos have focused on the contest within the state and markets, our work shows that concerns related to marked dynamics generalise to a broader design principle on the role of any kind of competition logistic in social media services.
Similarly, we observed differences in the formulation of user identity. Both left-leaning groups appeared to allow and encourage more flexible profile work, while the right-leaning group limited the options for users to present themselves. In part, this difference may speak to different conceptualisations of how people form social groups. For example, occupation as a primary social group is important for left-wing parties, which historically built on social class and related occupational status, know as class-mass parties (Gunther and Diamond, 2003). Therefore, it is not surprising that these fields for self-expression appeared more extensive. In addition, supporters of left-leaning parties tent to have more extraverted personality and be more open to experiences (Osborne et al., 2021). Therefore, it reasons to argue that they may embrace a more nuanced representation of identity.
@@ 5.2. Beyond Left and Right on the Ideological Spectrum
As highlighted in Section 2.1, left–right is only one ideology dimension. Our analysis and outcomes directly relate to our participant selection process, where we sought groups along this axis. We also collected users’ ideological positions on the liberal–conservative axis. Group 3 stands out as the only one with a pronounced liberal-leaning ideological measurement (see Table 1). Other groups had only a slight leaning towards the liberal end but had a high variance within the group of participants, indicating the groups were more heterogeneous on this axis.
That might explain why Group 3 was the only group that encouraged the idea of an algorithmically and profile-based structured social network and open network interactions. This form differs from the more traditional approach of structuring and interaction and might even signify an underlying agreement of slightly more conservative group identities (Lampinen et al., 2014). This points to a disparity between the two left-leaning groups and highlights the importance of investigating the influence of political ideologies in the design process of digital technologies beyond the left-right spectrum.
@@ 5.3. Limitations and Future Work
There are several limitations to our study: most vitally, our sample is small, and there is a lack of right-leaning and mixed groups, inhibiting a more extensive within-ideology comparison. We re-iterate that this was due to the challenge of reaching out to a particular political party, thus limiting our access to right-leaning participation. Nonetheless, the small sample may limit the generalisability of our findings. Similarly, we acknowledge that some results might be due to the difference in the groups’ social networking experiences, age, or other compounding factors. However, as we connect our findings to political ideologies, the results have a stronger foundation on political science than just providing empirical observations.
@@
Second, political ideologies are highly contextualised. The outcomes may be different in another country or may change temporarily as political ideologies and their thinking evolve (for summary, see Cochrane, 2015b). For example, Nelimarkka et al. (2019) show that acceptable polarisation mitigation strategies for social media differ in Finland and the United States. These two issues call for larger studies and comparative approaches to examine how the political ideologies shape potential design language.
@@
Finally, this analysis focused only on final artifacts and their presentation. Further research could examine more group dynamics during the group design stages. It would be possible to analysis to examine who introduced design ideas, how those were discussed and if those discussions lead to their incorporation in the final artifact. This would provide more nuanced perspectives on where the design ideas emerge and how they are examined in the group, thus also addressing power stances.
5.4. Implications to Practice
Our work emphasizes the importance of political ideologies as coherent value sets in the design of social media services. We begin to build a connection between the state – represented by its democratically elected political representatives affiliated with political parties – and technology design beyond mere regulation. Traditionally, political action in the technology domain has focused on regulating the relationship between users and platforms, such as through GDPR or the AI Act, and on regulating interactions between platform companies, such as in antitrust cases. However, we propose that there is a more general design language that outlines acceptable affordances in technology design per political ideology.
Therefore, researchers using value-based approaches or seeking to introduce ideas of ‘democratic design’ should ensure that diversity on the ideological spectrum is maintained to avoid political leanings in such works, especially if touching on applications with societal relevance. Furthermore, our work diversifies the focus on ideologies, whereas current critical scholarship focuses mainly on capitalism and feminism (Keyes et al., 2019; Bardzell, 2010). Our work highlights the diversity of ideologies and their utility in examining values.
We would not advocate a generic social media service to examine ideological diversity, as they operate across various settings. (However, we do consider it a relevant question which ideological design principles are in use in these services, cf., Kender and Frauenberger, 2022; Hasinoff and Bivens, 2021) However, for projects planning to expand social media to serve as a better avenue for political discussion, political ideals of participation are essential – concepts such as good democracy are ultimately normative (Held, 2006). Such efforts may forget how technologies manifest values, and choices such as self-presentation or competition for visibility are entangled with political values. In such cases, deliberative processes where a service is developed in a mixed ideology setting are vital to ensure the acceptability of the service across the ideological spectrum. Our mixed composition group already seemed to integrate both more left-leaning and right-leaning principles into their artifact.
@@ 6. Conclusions: Towards Political HCI?
Our study explored the design of social media services within groups formed by politically aligned elites. By examining the artifacts produced, we observed that these designs reflect ideological alignments: certain features are preferred in groups with right-leaning members, while others are favored by left-leaning members. Our findings underscore that political ideologies represent a coherent set of values even in the design domain of digital systems, extending and generalizing existing political science research on ideologies.
Political power increasingly influences the development of social computing systems through legislation and government funding for technology development. Beyond this direct influence, political ideologies manifest as broader sets of values, raising the challenge of ensuring more democratic representation of these values during system development. Therefore, social computing researchers must engage with political science and the study of political ideologies. This engagement will help understand, unpack, and make visible the political power embedded in these systems, fostering public debate and further examining digital systems’ societal impact.

### v2

@@ 5. Discussion
We began our work by highlighting a challenge related to how values are often considered in human–computer interaction without focus on political elites and their political ideologies. We show that political ideologies should be considered as a relevant set of values in the context of technology design.
Overall, there were some commonalities in the design approaches. For example, all groups identified the need for strong authentication, linking the users real identity with their online profile. Similarly, all groups included a picture in the user profile. We also observed differences across the groups. Some of these differences align with previous research on political ideologies and their characteristics in non-digital domains. This highlights the relevance to examine political ideologies as part of the design work.
@@ 5.1. Design by the Left and the Right
Previous work in political science has various characterisations of the economic left and the right: while somewhat unclear labels, both citizens and experts identify this continuum and can position themselves into it (see e.g., Cochrane, 2015b). Using political party manifestos, scholars such as Cochrane (2015a), Jahn (2011) and Däubler and Benoit (2022) have identified common characteristics of political parties based on their ideological leaning: right-leaning parties are characterised with lesser degree of support to state intervention in the economics or to more expansive welfare programs (see Table 5, first row). As these are from political party manifestos, these do not directly map into the design of digital services. Rather, they broadly characterise the division between the left and the right in their ways to approach societal questions.
@@
Moving from political manifestos to digital artefacts and design principles in them (see Table 5, second row). Grön and Nelimarkka (2020) suggested that the left–right dimension appeared as a tension between paternalism and freedom. More extensively, they showed acceptable design approaches seemed to be rooted in the values of each party, for example, communality was preferred by the liberal-conservative and agrarian Centre Party. Our work continues to examine this area, showing alignment of our work with the work on political manifestos and previous work on digital artefacts and ideologies.
We observed a difference in design choices regarding popularity-based visibilitythat can be interpreted to correspond group’s ideological compositions. The right-leaning group suggested affordances encouraging popularity as a driver for content visibility, which seems to point towards an underlying idea of the platform as a marketplace for ideas competing for attention. Both left-leaning groups explicitly discouraged such an approach, even criticising such approaches. This difference continues the well-established distinction of the economic left and right in the competitive logics: left-leaning parties do not in general prefer it, while right-leaning parties believe that open markets and competition leads to positive gains to the society (see Table 5, first row). However, we move this observation from the party manifestos and non-digital choices to the digital service design.
Similarly, we observed differences in the formulation of user identity. Both left-leaning groups appeared to allow and encourage more flexible profile work, while the right-leaning group and mix-membership group limited the options for users to present themselves. In part, this difference may speak to different conceptualisations of how people form social groups. For left-leaning parties, group identities have historically been important as attributes which unite people: they were organised as class-mass parties and therefore their support has followed social class and related occupational status (Gunther and Diamond, 2003). Still today, collective identities are important for the left; among other things as an approach to examine and organise around societal fairness (Bernstein, 2005). Therefore, it is reasonable to argue that they may embrace a more nuanced representation of identity.
Overall, we further the argument that political ideals of the left and right guide how digital topics are approached (see Table 5). We see that similar themes touched on non-digital domains, such as party manifestos, can also be identified in the design of digital artefacts and their affordances. This highlights how political ideologies may – at least partly – shape how participants perceive the problem and choose what the acceptable solution spaces are for such problems. Our findings on competitive logic and identity and profile work are aligned with the political manifesto work, but also show how profound aspects of social media design are connected with ideologies: these relate to well-established core affordances of social media services.
@@ 5.2. Beyond Left and Right on the Ideological Spectrum
As highlighted in Section 2.1, economic left–right is one ideological dimension; but there are other ideologies as well. Our analysis and outcomes directly relate to our participant selection process, where we sought groups along the economic left–right dimension for this study. However, following the two-dimensional model, we know that Group 3 stands out as the only one with a pronounced liberal-leaning ideological measurement (see Table 1). Other groups had only a slight leaning towards the liberal end, but had a high variance within the group of participants, indicating that the groups were more heterogeneous on this ideological dimension.
Among the observations not explained above was Group 3’s choice to focus on an algorithmically and profile-based structured social network that allows open network interactions. The “forum” solution by the other groups follows a more traditional approach of structuring the interaction in clearly defined subgroups, which limits network effects. This might even signify an underlying agreement of slightly more conservative values across party lines and a difference between the two left-leaning groups.
Given our small sample, our evidence is only an indicator and requires further research. Nonetheless, we make this remark as a response to how the HCI and CSCW research community engages political ideologies with only one dimension (e.g. as in Zhang and Counts, 2015b, p. 2607). While this may be acceptable within the United States, the narrow perspective on ideologies cannot be sustained in a global community – even Western Europe differs from this in the additional value dimension. In addition, political scientists have highlighted that the ideological landscape is currently changing (Ford and Jennings, 2020a; Jahn, 2023) , thus requiring sensitivity on this topic.
5.3. On the method: How to examine ideologies of artefacts?
It is often challenging to work out design reasoning from presentations and artefacts; as we note, the framework we used for analysis (Hasinoff and Bivens, 2021) was not specific on how to make this move. Similar challenges also exist for other works which seek to extract ideologies from artefacts (Light et al., 2018). In response to this challenge, we expanded the analysis approach and embarked on interpretative work on the affordances the participants had included in their designs by reflecting on them using our knowledge of the self-described ideological leanings of the designers as a lens for analysis.
Self-described ideological leanings provide us an additional detail for interpretative work; but in qualitative research there are position towards the interpretation (e.g., Soden et al., 2024; Braun and Clarke, 2019). Willig (2017) divides interpretation on a continuum between suspicious and emphatic approaches. Our approach falls towards the suspicious end of the continuum: the researchers work to connect and interpret observations to wider theoretical perspectives instead of focussing on how participants themselves provide the meanings; this latter is referred to as the emphatic approach. To ensure transparency in this work, we have used excerpts from the research material to justify our interpretation and used established works in political science to further situate our findings in established body of work.
An emphatic approach would have been possible as well: we could have asked the participants to directly explain their choices from each member’s political viewpoint. In an individual setting, Grön and Nelimarkka (2020) used such more emphatic approach to elicit an explicit connection with the ideology of the party and the party and the choice of design. However, as our study was a collaborative design work, we did not want to prime participants at any point of time to focus on their parties’ political viewpoints.
While a suspicious approach limits our outcomes (see next section), it allows us to examine how design work is discussed and when and how political alignments step into the discussion and design choices. Any overt priming to political parties may have lead the design workshop to become more a debate of party positions than to work towards design itself. At the same time, as the self-described ideological positions were known to researchers, we build a bridge from ideologies to artifacts.
@@ 5.4. Limitations and Future Work
There are several limitations to our study: most vitally, our sample is small, and there is a lack of right-leaning and mixed groups, inhibiting a more extensive within-ideology comparison. We re-iterate that this was due to the challenge of reaching out to a particular political party, thus limiting our access to right-leaning participation. Access is a known problem when studying elites (Jupp, 2006). Nonetheless, the small sample limits the generalisability of our findings. Similarly, we acknowledge that some results might be due to the difference in the groups’ social networking experiences, age, design experience, or other compounding factors. For example, it may be that more experienced users of social media could reflect on various existing design approaches and thus have wider approaches in their minds. Teasing these differences would require a much more extensive number of groups; previous studies on the impacts of political ideology on deliberation had 32 groups and 256 participants (Grönlund et al., 2015). However, while our sample is clearly underpowered to draw such conclusions, we connect our findings to previous work regarding political ideologies to give them a stronger foundation. Our results highlight that similar patterns observed in the non-digital domain translate as design inspirations in the digital domain.
@@
Second, political ideologies are highly contextualised. The outcomes may be different in another country or may change temporarily as political ideologies and their thinking evolve (for summary, see Cochrane, 2015b). For example, Nelimarkka et al. (2019) show that acceptable polarisation mitigation strategies for social media differ in Finland and the United States. These two issues call for more extensive studies and comparative approaches to examine how political ideologies shape potential design language, which is beyond the scope of this study design.
@@
Finally, this analysis focused only on the final artefacts and their presentation. Further research could examine group dynamics during the design stages. It would be possible to analyse who introduced design ideas, how those were discussed and if those discussions led to their incorporation in the final artefact, focussing on the interaction between participants to understand themes such as power and tensions (previously studied in more generic participatory design by Bratteteig and Wagner, 2012). We envision this direction as an area for future work to gain a nuanced perspective on how the design ideas emerge and then are shaped by the group into the final design outcomes examined in this work.
5.5. Implications to the Design of Social Media
The CSCW community has worked extensively on the design of social media to mitigate societal concerns such as political polarisation (e.g., Nelimarkka et al., 2018; Munson et al., 2009), derogatory language use and poor argumentation (Seering et al., 2019; Jia et al., 2024; Kriplean et al., 2012), or lack of civic control in these systems (Zhang et al., 2020; Fan and Zhang, 2020). Given these concerns, there are proposals on alternative value bases for social media services. Various works have examined how algorithmic changes (e.g., Jia et al., 2024; Combs et al., 2023; Feltwell et al., 2020; Garimella et al., 2017) or alternative design approaches (e.g. Semaan et al., 2014, 2015; Nelimarkka et al., 2018) can mitigate these concerns.
However, these approaches have not always gained political support due to ideological conflicts. The lack of political support is most prominently seen in the debate on moderation and free speech: Haapoja et al. (2020) elaborates how the deployment of a hate speech detection system produced a counter-movement from those who considered it too limiting for political expression; this divide aligned with the tensions between political ideologies. These examples show how the ideological divisions – similar to those observed in our study and Grön and Nelimarkka (2020) – have direct implications on how a service operates.
Therefore, the research community working to propose alternatives should diversify their ideological considerations to capture design space more comprehensively. For example, the recent proposal to order the news feed based on a societal objective function (Jia et al., 2024) instead of the engagement-based news feed omits that specifying such a function might be impossible in a pluralistic society. Their work chooses the function from political science work on anti-democratic behaviours. However, working the function out from political sciences does not make it free from values. For example, conceptualisations of democracy differ widely (Held, 2006), and thus what counts as anti-democratic depends ultimately on a choice about values.
When aiming to ‘fix’ social media’s downsides, we must acknowledge that even clearly ideologically motivated work could benefit from examining alternative ideological stances. Fixing efforts may overlook how technologies manifest values, as choices like free self-presentation or competition for visibility are entangled with established political values, as shown in this research. One potential direction is to form deliberative processes where a service is developed in a mixed ideology setting vital to ensure the acceptability of the service across the ideological spectrum. Our mixed composition group already seemed to integrate both more left-leaning and right-leaning principles into their artefact.
Alternatively, we may need to conclude that a service cannot be redesigned to cater for all perspectives. Each service has their users and context of use. For example, users of currently conservative Truth Social may find some design directions unacceptable overall, and might not find it relevant to incorporate more liberal perspectives to the design work at all. Similarly, cultural differences may impact acceptable and unacceptable design directions (Nelimarkka et al., 2019). Indeed, recent changes due to Elon Musk’s transformation of Twitter into X and related changes e.g., on moderation have sparked similar migration of users to other services (e.g. BlueSky or Mastodon). Hence, there is always the avenue to design new services to provide a space for those who feel not aligned with the platform. Furthermore, approaches such as federation would allow platforms share some areas of the content.
5.6. Implications to Politics and Design
Beyond the specifics of social media design, our work emphasises the importance of political ideologies as value sets, showing that the design approaches chosen can be interpreted as being rooted in broader ideological aims held by people. It has been understood that policy and politics are essential parts of design activity (e.g., Jackson et al., 2014; Kannabiran and Petersen, 2010) and there are various works which focus on designing from a particular ideological stance (e.g., Fox et al., 2019).
We expand this discussion by building a connection between the democratically elected political representatives affiliated with political parties and technology design beyond mere regulation. Traditionally, political action in the technology domain has focused on regulating the relationship between users and platforms, such as through GDPR or the AI Act, and controlling cooperation between platform companies, such as in antitrust cases. Also here, political ideologies are at the core of these actions. Political ideologies shape what goes into regulations and laws, and the digital is increasingly an area for politics. Indeed, political party manifestos clearly differ on how they address the democratic potential of digital platforms, and organisations of labour and capital relationships differ (Guglielmo, 2024). Guglielmo (2024) argues that these create new lines of tensions along which parties differentiate each other. Beyond these ideas of how the digital is seen, our work indicates a more general design language that outlines acceptable affordances in technology design per political ideology.
Therefore, our conclusion from the prior section for a need to diversify ideological perspectives in the design and research process applies not just to social media but more generally to design, particularly to areas where societal implications are clear, and more extensively to value-sensitive and critical design approaches. Current the academic discussion on ideologies has mainly focused on two streams: capitalism-critique and feminism (e.g., Keyes et al., 2019; Bardzell, 2010) or even see ideologies as “constraints” to be resolved (Burrell and Metcalf, 2024). We call for expanding how we approach and acknowledging ideologies: ideologies should be explicit, both in the data and author’s position.
This may require asking participants to identify their ideological stance, either directly using dimensions like economic left–right, or through value statements, and reporting these as background factors similar to commonly used age, gender, and race/ethnic background; and ensuring that ideological leanings are transparent. Additionally should be reflective about how their values are placed in the spectrum of ideologies. Similarly, author positionality, particularly on various critical design approaches, such engage more deeply with ideological commitments and if possible, reflect how other kinds of commitments may manifest themselves.
Underlining both of these observations, we have used the conceptualisation of political ideologies as a springboard for more extensive engagement with politics and policy. As our study indicates, this conceptualisation and its relationship with governments and politics may provide a fruitful angle to attack the policy–design gap.
@@ 6. Conclusions: Towards Political HCI?
Our study explored the design of social media services with four groups formed by politically aligned elites. By examining the artifacts produced, we observed that some of the design choices were aligned with elite’s ideologies. Certain features were preferred in groups with right-leaning members, while other features were favoured by left-leaning members. Our findings underscore that political ideologies represent a coherent set of values even in the design domain of digital systems, extending and generalizing existing political science research on ideologies.
Political power increasingly influences the development of information technology and social computing systems through legislation and government funding for technology development. Beyond this direct influence, political ideologies manifest as broader sets of values, raising the challenge of ensuring more democratic representation of these values during system development. Therefore, social computing researchers must engage with political science and the study of political ideologies. This engagement will help understand, unpack, and make visible the political power embedded in these systems, fostering public debate and further examining digital systems’ societal impact.